OpenStudy (marcelie):

evaluate the integral :

OpenStudy (marcelie):

\[\int\limits_{}^{} \frac{ \sec^6(x) }{ \tan^2(x)}\]

OpenStudy (marcelie):

do i convert the sec to 1/cos^2(x) ?

Nnesha (nnesha):

hmm have to use this identity 1+tan^2x =sec^2x \[\huge\rm \int\limits_{ }^{ } \frac{ \frac{1}{\cos^6x }}{ \frac{\sin^2x}{\cos^2x} } \] \[\int\limits_{ }^{ } \frac{1}{\cos^6x} \cdot \frac{\cos^2x}{\sin^2x}\] this looks mssy hmm

Nnesha (nnesha):

sec=1/cosx btw

OpenStudy (marcelie):

can we cross out the 2 signs and we will be left with \[\frac{ 1 }{ \cos^4(x) *\sin^2(x) }\]

OpenStudy (legomyego180):

rewrite it \[\frac{ \sec^2(x)\sec^2(x)\sec^2(x) }{ \tan^2(x)}\] Remember the identity: \[\tan^2(x)=\sec^2(x)-1\]

Nnesha (nnesha):

that's not gonna work bec if you use sin^2x=(1-cos(2x))/2 identity it would not help you to cancel cos(x) or let u=cos x....

OpenStudy (legomyego180):

https://www.youtube.com/watch?v=xFrO9gUYdB4

OpenStudy (marcelie):

im lost @_@

Nnesha (nnesha):

\[\int\limits_{ }^{ } \frac{\sec^6x}{\tan^2x}\] we can use this identity 1+tan^2x=sec^2 but first we need to rewrite to get sec^2x or tanx \[\int\limits_{ }^{ } \frac{\sec^6x}{ tanx \cdot tanx}\] if i split tan^2x i wouldn't be apply to use that identity therefore \[\frac{ \sec^4x \cdot \sec^2x }{ \tan^2x }\] or just like legomyego mentioned \[\frac{ \sec^2x \cdot \sec^2x \cdot sec^2x }{ \tan^2x }\] sec^2x=tan^2x +1 \[\int\limits_{ }^{} \frac{ (1+\tan^2x)(1+\tan^2x) \cdot \sec^2x }{ \tan^2x }\] let u =tanx hope that would work hmm

satellite73 (satellite73):

alternatively , you can make this \[\int(\tan^2(x)+1)^2(\cot^2(x)+1)dx\] although how you would recognize this is anyone's guess

OpenStudy (marcelie):

u sub ?

OpenStudy (sshayer):

\[I=\int\limits \frac{ \sec ^4x \sec ^2x }{ \tan ^2x }dx=\int\limits \frac{ \left( 1+\tan ^2x \right)^2\sec ^2x }{ \tan ^2x }dx\] \[=\int\limits \frac{ 1-2\tan ^2x+\tan ^4x }{ \tan ^2x }\sec ^2x dx\] put tan x=t \[\sec ^2x dx=dt\] \[I=\int\limits \frac{ 1-2t^2+t^4 }{ t^2 }dt\] ?

OpenStudy (marcelie):

hmm i kind of did what satelite did but i think it makes it more complicated ? or is there an easier way in doing these

OpenStudy (sshayer):

\[I=\int\limits t ^{-2 }dt-2\int\limits dt+\int\limits t^2 dt+c\]

Nnesha (nnesha):

that is the easy one when i see sec^nx and tan^mx i think abt 1+tan^2x=sec^2x :D

OpenStudy (sshayer):

correction use + 2tan^2 x in place of -2 tan^2x

OpenStudy (legomyego180):

@marcelie this is trig powers, its one of the hardest things you have to do in calc 2. Theres no easy way around them really. Luckily they dont really show up in any other class I hear except this one.

OpenStudy (marcelie):

lool yeah but are there like rules for these like when sin is odd and cos is odd for their powers?

OpenStudy (sshayer):

\[I=\int\limits \frac{ \left( 1+2 \tan ^2x+\tan ^4x \right) }{ \tan ^2x }\sec ^2x dx\]

Nnesha (nnesha):

\[\int\limits_{ }^{} \sin^ax ~~\cos^b x \] when a or b = odd apply this identity \[\sin^2x= 1-\cos^2x ~~~or ~~\cos^2x= 1-\sin^2x\] if a or b = even \[\sin^2(x) =\frac{ 1-\cos(2x) }{ 2}~~~~\cos^2(x)=\frac{1+\cos(2x)}{2}\]

OpenStudy (sshayer):

in the end \[I=\int\limits t ^{-2}dt+2 \int\limits dt+\int\limits t^2 dt+c\]

Nnesha (nnesha):

i think it would be easy if you sub u for tan\[\int\limits_{ }^{} \frac{ (1+u^2)(1+u^2) }{ u^2 } \sec^2x\]

OpenStudy (marcelie):

= \[(1+2u^2+u^4)/ u^2\]

OpenStudy (marcelie):

\[(u ^{-2})(1+2u^2+u^4)\]

Nnesha (nnesha):

looks good \[\int\limits_{ }^{} \frac{ 1+2u^2+u^4 }{ u^2 } \cancel{\sec^2x }\cdot \frac{du}{\cancel{\sec^2}}\] \[\int\limits_{ }^{} \frac{1}{u^2}+\frac{2u^2}{u^2}+\frac{u^4}{u^2} du \]

OpenStudy (marcelie):

\[\frac{ 1 }{ u^2 }+2+u^2\]

Nnesha (nnesha):

yes^^

Nnesha (nnesha):

now you can rewrite 1/u^2 = u^{-2} would be easy to integrate :D

OpenStudy (marcelie):

okay so then i got this \[-\frac{ 1 }{ u }+2u+\frac{ 1 }{ 3 }u^3\]

Nnesha (nnesha):

looks good and +Chocolates :)

OpenStudy (marcelie):

lool okie so then final answer: \[\frac{ 1 }{ 3 }\tan^3(x)+2\tan(x)-\frac{ 1 }{ \tan(x) }+c\]

OpenStudy (marcelie):

so that 1/tanx converts to cot (x)

Nnesha (nnesha):

hmm you don't have to but ofc u can .... i would leave it like that

OpenStudy (marcelie):

oh okayy :D

Nnesha (nnesha):

i love these type of questions <3 funn!! :D

OpenStudy (marcelie):

loool D: but question are there rules how to format these ? when its sec or tan

Nnesha (nnesha):

hmm use this identity 1+tan^2x=sec^2x doesn't matter if the exponents are odd or even but first rewrite the function to get sec^2x `or` tanx i always let u equal to base of highest exponent here is an example \[\large\rm \int\limits_{ }^{ } \sec^5x \cdot \tan^3x ~dx\] i will split tan^3x = tan^2x * tanx \[\large\rm \int\limits_{ }^{ } \sec^5x \cdot \tan^2x \cdot \color{REd}{tan x} ~dx\] apply the identity \[\large\rm \int\limits_{ }^{ } \sec^5x \cdot (sec^2x-1 ) \cdot \color{REd}{tan x} ~dx\] now let u = secx the reason why we need tan x or sec^2x because when we let u= something the derivative of that *something* should already in the function \[\large\rm \int\limits_{ }^{ } \sec^5x \cdot (1+sec^2x) \cdot \color{REd}{tan x} ~dx\] u= secx ( base of highest exponent ) du = secx tan x dx dx=du/(secx tanx) \[\int\limits_{ }^{ }\color{REd}{u^5}(u^2 -1) \tan x \cdot \frac{du}{secx ~tanx}\] u=sec u^5 is same as u^2 secx \[\int\limits_{ }^{ }\color{blue}{ u^4}(u^2 -1) \tan x \cdot \color{REd}{secx} \cdot \frac{du}{secx ~tanx}\]

Nnesha (nnesha):

not good at explaining stuff but hopefully that makes sense

OpenStudy (marcelie):

oh okay ill look at it right now

Nnesha (nnesha):

correction u=sec u^5 is same as `u^2 secx ` typoo it should be u^4 secx